OpenStudy (kabbledoo):

Let f(x) = x^2 − 81. Find f^−1(x).

OpenStudy (kabbledoo):

PLeaseee helpp

OpenStudy (photon336):

so this question is to find the inverse of this function right

OpenStudy (photon336):

first switch x and y

OpenStudy (kabbledoo):

i think so

OpenStudy (photon336):

\[x = y^{2}-81 \]

OpenStudy (photon336):

so then you would solve for y to find the inverse

OpenStudy (kabbledoo):

im confused on that part

OpenStudy (photon336):

you get what I did in the first part right?

OpenStudy (kabbledoo):

yes

OpenStudy (photon336):

so now we have this \[x = y^{2}-81 \]

OpenStudy (photon336):

now to solve for y it's easy we have to get y by itself. \[x+81 = y^{2}\]

OpenStudy (photon336):

so far so good?

OpenStudy (kabbledoo):

okay thnk so

OpenStudy (photon336):

now we take the square root of both sides: \[\sqrt{x+81} = y \]

OpenStudy (kabbledoo):

oh so it would be plus or miuns?

OpenStudy (photon336):

I think that would be it

OpenStudy (kabbledoo):

yay thanks man